OpenStudy (anonymous):

Find f(x)^-1=sinx I know its cosx, but how do I get there?

OpenStudy (agent47):

y=Sinx inverse means: x=Siny, solve for y here and you'll get f^-1

OpenStudy (agent47):

y=Sin^-1(x)

OpenStudy (zzr0ck3r):

the inverse of sinx is the inverse of sinx, it is not cos(x) the only other way we rwite it is arcsin(x), so I am not quite sure what the question is

OpenStudy (anonymous):

Oh I see, I need to find the prime of y=sin(X), not inverse

OpenStudy (anonymous):

Its from khans video on L Hopital's rule. http://www.youtube.com/watch?v=PdSzruR5OeE @ 6:55

OpenStudy (anonymous):

Well I guess I owe you for helping me out earlier zzr0ck3r. I found out you use the quotient rule to solve this so time to find out how. Thanks!

OpenStudy (zzr0ck3r):

you cant solve for f^(-1)= sin(x) this is a function its like saying solve f(x) = 3x...

OpenStudy (zzr0ck3r):

O sorry I did not read your reply about the prime thing.. sorry I dont know what he means by prime.

OpenStudy (zzr0ck3r):

are you learning l'hopitals rule in class?

OpenStudy (anonymous):

No im learning l'hopitals rule for homework :(

OpenStudy (zzr0ck3r):

well, if you did not learn it in class, your teacher does not want you to use it for homework.

OpenStudy (zzr0ck3r):

or if you have not learned it in the chapter your working in.

OpenStudy (zzr0ck3r):

have you learned how to differntiate yet?

OpenStudy (zzr0ck3r):

o prime you mean derivative...

OpenStudy (anonymous):

Well its AP Calc and our teacher just expects us to know everything.

OpenStudy (zzr0ck3r):

have you learned how to differentiate?

OpenStudy (zzr0ck3r):

do you know what this is dy/dx = x^2?

OpenStudy (anonymous):

No i dont :(.

OpenStudy (zzr0ck3r):

if you have not then do not try and understand l'hopitals rule yet. this is what calc is about...you are trying to learn how to run a 5k race before you know how to crawl:)

OpenStudy (zzr0ck3r):

there is much needed to learn about limits before you can learn to differentiate, and you cant learn l'hopitals law if you dont know how to differentiate.

OpenStudy (anonymous):

Guess its time for me to hit the books.

OpenStudy (anonymous):

Time for me to go to bed Thanks for all the help today.

OpenStudy (zzr0ck3r):

stick with limits, that person should not have told you to use that rule, as you were saposed to show it with algebra. not calc.

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

Oh ok, Ill be sure to look surprised when we learn about the rule.